The complete combustion of 0.731 g of a snack bar in a calorimeter (Ccal = 6.15 kJ/°C) raises the temperature of the calorimeter by 1.93 °C. Calculate the food value (in Cal/g) for the snack bar.
1
Expert's answer
2020-03-09T09:45:36-0400

Here, it is given that mass of snack bar as 0.731 g ,

Temp of calorimeter is given as "\Delta T= 1.93 ^\circ"

for calorimeter "C" cal= 6.15 "\frac{ kJ}{^oC}"

So as per question , Heat energy released by the calorimeter = C "\Delta T"

Q= 6.15 "\times" 1.93 = 11.8695 kJ

heat energy is calorie= "\frac{11.8695 kJ}{4.184 \frac{kJ}{kcal}}"

Q= 2.8368 kcal

now as per question the food value for snack bar = "\frac{ Q}{mass of snack bar }"

food value = "\frac{2.8368 kcal }{0.731g} = \frac{2.8368 \times1000 cal}{0.731g}"

food value = 3880.8 "\frac{cal}{g}"

This is the required food value of snack bar